OpenStudy (anonymous):

solve a rectangular garden is three times as long as it is wide. if the area of the garden is 2523ft^2, find the length and width of the garden

hero (hero):

3w*w = 2523 3w^2 = 2523 w^2 = 841 w = 29 3w = 29*3